# The Quest for the Origins of Ultra-High-Energy Cosmic Rays

### v1

@@ 3 UHECR propagation
The propagation of UHECRs through intergalactic space is a complex process that depends on the properties of the particles themselves, the magnetic fields they traverse, and the background radiation fields with which they interact (see §3.1 and §3.2). Accurately modeling UHECRs propagation requires accounting for these factors, which are typically combined within computational tools designed to simulate their transport (see §3.3).
@@ 3.1 Interactions with matter and radiation
The interaction of UHECRs with matter is generally considered negligible due to the low density of gas in the intergalactic medium (IGM). However, it can be important in other environments (Owen et al., , 2018), playing a particularly special role in gamma-ray and neutrino emission by galaxy clusters (Fang and Olinto, , 2016; Hussain et al., , 2021, 2023).
@@ 3.2 Magnetic fields
The presence of extragalactic magnetic fields can significantly influence the propagation of UHECRs. Naturally, the EGMF between sources and Earth determines by how much CRs are deflected and how much longer they must travel before reaching Earth. Since most of the universe’s volume consists of “empty” cosmic voids, it is reasonable to expect that the fields in these regions dominate the propagation of UHECRs, although this remains to be proven. This is captured by the notion of volume filling factor, illustrated in fig. 3. One could, of course, argue that outflows from large-scale structures like galaxy clusters can inject magnetised material into the IGM, but this effect should not significantly alter the magnetisation of cosmic voids over the age of the universe (e.g., Furlanetto and Loeb, , 2001; Barai et al., , 2008; Beck et al., , 2013; Arámburo-García et al., , 2021). Therefore, unless sources are located very close to Earth such that they do not traverse cosmic voids, UHECRs would likely cross these regions at least once.
@@ 4.1 Centaurus A
  • If acceleration is electromagnetic, the size of the confining magnetised region must enable the escape of CRs with the energy of interest.
@@ 4.2 The starburst correlation
What is more: even if we can unambiguously confirm that UHECRs come from SBG, we are still not able to identify the acceleration mechanism. As described in §2, acceleration mechanisms are simply templates and can be applied to a variety of types of objects which could also contribute to the observed flux. Then the question becomes: what can we learn from the UHECRSBG correlation? On its own, at present, we have not learnt anything fundamentally new. However, this signal is a first step toward unveiling the sources of UHECRs. With higher statistics and better knowledge of propagation ingredients such magnetic fields and background photon fields, it is possible to add information from other messengers, in particular neutrinos and gamma rays, to build a self-consistent model that describes UHECR emission by SBG. Note that multi-messenger and multi-wavelength information are necessary conditions to meaningfully interpret these observations but they are still not sufficient.
@@
    large-scale phenomena inherent to SBG themselves, such as super-winds, are the accelerators;
@@ 4.3 Combined spectrum-composition fits
where $\kappa=\{\left(A,Z\right)\;|\;\left(A>0,\,Z\geq 0\right)\}$ is the set of all CR species emitted, with atomic number $Z$ and atomic mass $A$. Here $E$ is the energy of the emitted CR, $\mathcal{S}_{\varkappa}$ is a function that describes the spectrum, with a characteristic composition-dependent maximal energy ($E_{\text{cut}}^{(\varkappa)}$), and $t$ is a time variable, which is related to the redshift.
@@
Recently, Ehlert et al., (2023) adopted a phenomenological parametrisation for the cut-off, $E_{\text{cut}}^{(\varkappa)}=E_{\text{cut}}\left(A,Z\right)=E_{0}Z^{a}A^{b}$, wherein $E_{0}$ is a reference energy. In this case, Peters’ cycle (Peters, , 1961) is a special case of this parameterisation, with $a=1$ and $b=0$, and models with in-source photodisintegration (Unger et al., , 2015) have $a=0$ and $b=1$. They have shown that the $\mathcal{G}$ plays an important role in the outcome of these fits.
$\triangleright$ Source distribution. A common assumption in combined fits is that the UHECR sources are homogeneously distributed in space. This is a reasonable assumption for UHECRs of energies $E\lesssim 10\;\text{EeV}$, whose typical energy loss lengths are comparable to or larger than the distance to the sources, such that most sources are contributing to the observed flux. However, at higher energies ($E\gtrsim 40\;\text{EeV}$), the sources are mostly local (see fig. 2), lying within the interaction horizon ($\lesssim 100\;\text{Mpc}$), which means that they are not homogeneously distributed. More recently, some progress has been made in this direction, considering specific source populations such as AGN and minding the potential importance of possible nearby sources like Cen A to the observed flux (Pierre Auger Collaboration, 2024a, ). Moreover, new analyses looking for correlations with other messengers have also been informing us on the distribution of sources (Partenheimer et al., , 2024).
@@
$\triangleright$ Extragalactic magnetic fields. Most works, for simplicity, ignore the effects of EGMFs. This could be justified, but at present we lack information to support it (see §3.2). Wittkowski for the Pierre Auger Collaboration, (2017) performed a combined spectrum-composition fit considering a “realistic” distribution of EGMFs. They found that the inclusion of EGMFs could lead to a significant change in the results of the fits. For instance, in one of their scenarios, they inferred a spectral index of $\alpha\approx 1.61$ and a maximal rigidity of $\log(R_{\text{cut}}/\text{eV})\approx 18.88$, compared to $\alpha\approx 0.61$ and $\log(R_{\text{cut}}/\text{eV})\approx 18.48$, in the case where EGMFs were ignored. More recently, the Pierre Auger Collaboration, 2024b performed a similar fit considering the effects of turbulent EGMFs statistically (with one-dimensional simulations), which would imply a magnetic horizon effect on the UHECR spectrum. Even though this model is completely unrealistic, considering present-day knowledge of EGMFs (for reviews, see e.g., Alves Batista and Saveliev, , 2021; Vachaspati, , 2021), it is a step forward compared to most studies which assume an unmagnetised universe. And most importantly, it does confirm the aforementioned trend that the best-fit spectral indices tend to increase (for an $E^{-\alpha}$ spectrum). This has important implications for understanding the transition between Galactic and extragalactic CRs (Aloisio et al., , 2012; Parizot, , 2014), as magnetic horizon effects determine where the transition occurs (see, e.g., Alves Batista and Sigl, , 2014, for a discussion).
@@
Recently, the Pierre Auger Collaboration, 2024a extended the combined spectrum-composition fit by including an additional observable: anisotropies. Nevertheless, the model does not properly account for the effects of EGMFs, modelling it as a directional smear around the sources. UHECR trajectories are elongated due to magnetic deflections (see eq. 4), which raises interaction rates during propagation and can substantially affect the observed spectrum, composition, and anisotropies. Previously, Eichmann et al., (2018) had already performed a combined fit of the three observables (spectrum, composition, and arrival directions), focusing on radio galaxies. Such multi-observable fits are essential and should become the norm, given their constraining power. Yet, they inevitably open a “Pandora’s box” of modelling, drastically increasing the number of free parameters and hence complexity of the model.
@@ 5 Conclusions and Outlook
We find ourselves at a crossroads of UHECR astrophysics. Existing facilities such as the Pierre Auger Observatory and the Telescope Array are expected to continue operating for another decade, but it is unclear what will happen next (Coleman et al., , 2023). Plans are underway for new dedicated UHECR observatories like the Global Cosmic-ray Observatory (GCOS(Hörandel for the GCOS Collaboration, , 2022; Alves Batista for the GCOS Collaboration, , 2023) and POEMMA (POEMMA Collaboration, , 2021), in addition to observatories primarily designed for other messengers but which can also deliver information about UHECRs, such as the Giant Radio Array for Neutrino Detection (GRAND Collaboration, , 2020).
@@
Even if these new observatories are built, interpreting UHECR observations is far from trivial. Major theoretical challenges persist, and the assumptions made in interpreting the data are often unjustified and at times manifestly unrealistic. Understandably, this could be useful to make a problem tractable for a first attempt at tackling it, but it should not become the standard approach. Instead, models that progressively incorporate additional layers of complexity, besides being more robust, are more likely to provide a more accurate picture of the UHE universe.
Here I selected examples that illustrate why current models must be improved to provide a meaningful interpretation of the data, and how this can be done. The phenomenological fits described in §4.3, for example, can be promptly enhanced with existing tools, although the vastness of the parameter space of uncertainties still prevents an exhaustive scan. Still, some progress can be made without awaiting the next generation of observatories. The observations discussed in §4.1 and §4.2, on the other hand, do require next-generation instruments to enable actual discoveries.
@@ Acknowledgements
I acknowledge support from the Agence Nationale de la Recherche (ANR), project ANR-23-CPJ1-0103-01. I also thank the organisers and participants of ECRS2024, who contributed to the discussions that inspired this work.

### v2

@@ 3 UHECR propagation
The propagation of UHECRs through intergalactic space is a complex process that depends on the properties of the particles themselves, the magnetic fields they traverse, and the background radiation fields with which they interact (see §3.1 and §3.2). Accurately modelling UHECR propagation requires accounting for these factors, which are typically combined within computational tools designed to simulate their transport (see §3.3).
@@ 3.1 Interactions with matter and radiation
The interaction of UHECRs with matter is generally considered negligible due to the low density of gas in the intergalactic medium (IGM). However, it can be important in other environments (Owen et al., , 2018), playing a particularly special role in gamma-ray and neutrino emission by galaxy clusters (Fang and Olinto, , 2016; Hussain et al., , 2021; Condorelli et al., , 2023; Hussain et al., , 2023).
@@ 3.2 Magnetic fields
The presence of extragalactic magnetic fields can significantly influence the propagation of UHECRs. Naturally, the EGMF between the sources and the Earth determines by how much CRs are deflected and how much longer they must travel before reaching Earth. Since most of the universe’s volume consists of “empty” cosmic voids, it is reasonable to expect that the fields in these regions dominate the propagation of UHECRs, although this remains to be proven. This is captured by the notion of volume filling factor, illustrated in fig. 3. One could, of course, argue that outflows from large-scale structures like galaxy clusters can inject magnetised material into the IGM, but this effect should not significantly alter the magnetisation of cosmic voids over the age of the universe (e.g., Furlanetto and Loeb, , 2001; Barai et al., , 2008; Beck et al., , 2013; Arámburo-García et al., , 2021). Therefore, unless sources are located very close to Earth such that they do not traverse cosmic voids, UHECRs would likely cross these regions at least once.
In addition to voids, other regions of the large-scale structure of the universe also influence the propagation of UHECRs, affecting their trajectories. For instance, these structures can act as scattering centres (Kotera & Lemoine, , 2008), potentially causing significant deflections. They can also trap CRs particularly when sources are embedded within them (Alves Batista et al., , 2017; Hackstein et al., , 2018).
@@ 4.1 Centaurus A
  • If acceleration is electromagnetic, the size of the confining magnetised region must enable the escape of CRs with the energies of interest.
@@ 4.2 The starburst correlation
What is more: even if we can unambiguously confirm that UHECRs come from SBG, we are still not able to identify the acceleration mechanism. As described in §2, acceleration mechanisms are simply templates and can be applied to a variety of types of objects which could also contribute to the observed flux. Then the question becomes: what can we learn from the UHECRSBG correlation? On its own, at present, we have not learnt anything fundamentally new. However, this signal is a first step toward unveiling the sources of UHECRs. With higher statistics and better knowledge of propagation ingredients such as magnetic fields and background photon fields, it is possible to add information from other messengers, in particular neutrinos and gamma rays, to build a self-consistent model that describes UHECR emission by SBG. Note that multi-messenger and multi-wavelength information are necessary conditions to meaningfully interpret these observations but they are still not sufficient.
@@
    large-scale phenomena inherent to SBG themselves, such as super-winds, acting as accelerators;
@@ 4.3 Combined spectrum-composition fits
where $\kappa=\{\left(A,Z\right)\;|\;\left(A>0,\,Z\geq 0\right)\}$ is the set of all CR species emitted, with atomic number $Z$ and atomic mass $A$. Here $E$ is the energy of the emitted CR, $\mathcal{S}_{\varkappa}$ is a function that describes the spectrum, with a characteristic composition-dependent maximal energy ($E_{\text{cut}}^{(\varkappa)}$), and $t$ is a time variable, which may be related to the redshift ($t=t(z)$) or to the intrinsic variability of the source.
@@
Recently, Muzio et al., (2024) adopted a phenomenological parametrisation for the cut-off, $E_{\text{cut}}^{(\varkappa)}=E_{\text{cut}}\left(A,Z\right)=E_{0}Z^{a}A^{b}$, wherein $E_{0}$ is a reference energy. In this case, the Peters’ cycle (Peters, , 1961) is a special case with $a=1$ and $b=0$, and models with in-source photodisintegration (Unger et al., , 2015) have $a=0$ and $b=1$. They have shown that the $\mathcal{G}$ plays an important role in the outcome of these fits. Moreover, variations of $\mathcal{G}$ across a source population can also play a role (Ehlert et al., , 2023).
$\triangleright$ Source distribution. A common assumption in combined fits is that the UHECR sources are homogeneously distributed in space. This is a reasonable assumption for UHECRs of energies $E\lesssim 10\;\text{EeV}$, whose typical energy loss lengths are comparable to or larger than the distance to the sources, such that most sources are contributing to the observed flux. However, at higher energies ($E\gtrsim 40\;\text{EeV}$), the sources are mostly local (see fig. 2), lying within the interaction horizon ($\lesssim 100\;\text{Mpc}$), which means that they are not homogeneously distributed. More recently, some progress has been made in this direction, considering specific source populations such as AGN and minding the potential importance of possible nearby sources like Cen A to the observed flux (Pierre Auger Collaboration, 2024a, ). Moreover, new analyses looking for correlations with other messengers have also been informing on the distribution of sources (Partenheimer et al., , 2024).
@@
$\triangleright$ Extragalactic magnetic fields. Most works, for simplicity, ignore the effects of EGMFs. This could be justified, but at present information is lacking to support it (see §3.2). Wittkowski for the Pierre Auger Collaboration, (2017) performed a combined spectrum-composition fit considering a “realistic” distribution of EGMFs. They found that the inclusion of EGMFs could lead to a significant change in the results of the fits. For instance, in one of their scenarios, they inferred a spectral index of $\alpha\approx 1.61$ and a maximal rigidity of $\log(R_{\text{cut}}/\text{eV})\approx 18.88$, compared to $\alpha\approx 0.61$ and $\log(R_{\text{cut}}/\text{eV})\approx 18.48$, in the case where EGMFs were ignored. More recently, the Pierre Auger Collaboration, 2024b performed a similar fit considering the effects of turbulent EGMFs statistically (with one-dimensional simulations), which would imply a magnetic horizon effect on the UHECR spectrum. Even though this model is completely unrealistic, considering present-day knowledge of EGMFs (for reviews, see e.g., Alves Batista and Saveliev, , 2021; Vachaspati, , 2021), it is a step forward compared to most studies which assume an unmagnetised universe. And most importantly, it does confirm the aforementioned trend that the best-fit spectral indices tend to increase (for an $E^{-\alpha}$ spectrum). This has important implications for understanding the transition between Galactic and extragalactic CRs (Aloisio et al., , 2012; Parizot, , 2014), as magnetic horizon effects determine where the transition occurs (see, e.g., Alves Batista and Sigl, , 2014, for a discussion).
@@
Recently, the Pierre Auger Collaboration, 2024a extended the combined spectrum-composition fit by adding another observable: anisotropies. Nevertheless, the model does not properly account for the effects of EGMFs, simplifying it to a simple smear around the sources. UHECR trajectories are elongated due to magnetic deflections (see eq. 4), which raises interaction rates during propagation and can substantially affect the observed spectrum, composition, and anisotropies. Previously, Eichmann et al., (2018) had already performed a combined fit of the three observables (spectrum, composition, and arrival directions), focusing on radio galaxies. Such multi-observable fits are essential and should become the norm, given their constraining power. Yet, they inevitably open a “Pandora’s box” of modelling, drastically increasing the number of free parameters and hence complexity of the model.
@@ 5 Conclusions and Outlook
We find ourselves at a crossroads of UHECR astrophysics. Existing facilities such as the Pierre Auger Observatory and the Telescope Array are expected to continue operating for another decade, but it is unclear what will happen next (Coleman et al., , 2023). Plans are underway for new dedicated UHECR observatories like the Global Cosmic-ray Observatory (GCOS(Hörandel for the GCOS Collaboration, , 2022; Alves Batista for the GCOS Collaborators, , 2023; Ahlers et al., , 2025) and POEMMA (POEMMA Collaboration, , 2021), in addition to observatories primarily designed for other messengers but which can also deliver information about UHECRs, such as the Giant Radio Array for Neutrino Detection (GRAND Collaboration, , 2020).
@@
Even if these new observatories are built, interpreting UHECR observations is far from trivial. Major theoretical challenges persist, and the assumptions made in interpreting the data are often unjustified and at times manifestly unrealistic. Understandably, this can be useful for making a problem more tractable in an initial attempt to tackle it, but it should not become the default approach. Instead, models that progressively incorporate additional layers of complexity are not only more robust, but also more likely to provide an accurate representation of the UHE universe.
Here I selected examples that illustrate why current models must be improved to provide a meaningful interpretation of the data, and how this can be done. The phenomenological fits described in §4.3, for example, can be promptly enhanced with existing tools. Howerver, the vast parameter space of uncertainties still prevents a comprehensive exploration. Even so, some progress can be made without waiting for the next generation of observatories. The observations discussed in §4.1 and §4.2, on the other hand, do require next-generation instruments to enable actual discoveries.
@@ Acknowledgements
I acknowledge support from the Agence Nationale de la Recherche (ANR), project ANR-23-CPJ1-0103-01. I thank Michael Unger for useful comments on this manuscripts. I also thank the organisers and participants of ECRS2024, who contributed to the discussions that inspired this work.